OpenStudy (anonymous):

find a function f that satisfies the condition f''(x) = x^2, f'(0) = , f(0) = 3 this is antiderivatives and indefinite integrals (ie. applications of it)

OpenStudy (anonymous):

f''(x)=x^2 implies that f(x)= x^4/12+Ax+B f(0)=3 implies that B=3 f'(0)= gap implies A=gap so the function is f(x)=x^4/12 +gap x+3

OpenStudy (anonymous):

:S huh?

OpenStudy (anonymous):

I guess you missed out something?

OpenStudy (anonymous):

f'(0) =

OpenStudy (anonymous):

6

OpenStudy (anonymous):

ok than gap=6 :-)

OpenStudy (anonymous):

can u write it out w/ out the gap because i still dont see it

OpenStudy (anonymous):

f''(x)=x^2 implies that f(x)= x^4/12+Ax+B f(0)=3 implies that B=3 f'(0)=6 implies A=6 so the function is f(x)=x^4/12 +6x+3

OpenStudy (anonymous):

if something is not clear just ask

OpenStudy (anonymous):

oh ok i get it now lol...thank you sorry my eyes are so blurry